# Asynchronous mass inversion enriched quantum anomalous Hall states in multilayer graphene

### v1

@@ Abstract
Figure 3: (a). Side view of the Bernal tetralayer graphene (BTG), the blue and red atoms represent the $A$ and $B$ sublattices, respectively. The hopping parameters are indicated by different colors ($v_{i}=\frac{\sqrt{3}}{2}\gamma_{i}$): $\gamma_{0}=3000~{}meV$, $\gamma_{1}=400~{}meV$, $\gamma_{2}=-20~{}meV$, $\gamma_{3}=300~{}meV$, and $\gamma_{4}=\gamma_{5}=40~{}meV$. Here, the displacement field and staggered layer orders are set to zero: $\Delta=0$, and $m=0$. (b). The band structure near $K$ (or $K^{\prime}$) valley in the presence of the trigonal warping. The bands in the green circle show the enlarged view near $K$ (or $K^{\prime}$) point. (c). The band structure for the low-energy effective model along $k_{y}=0$ line with and without $\gamma_{3}$, which are shown by red solid and blue dashed lines, respectively. Here, we set $\Delta=0$ and $m=0$. (d). Enlarged view of the band structure for the region marked by the green circle in (c).
and the staggered layer order is written as:
@@
where $U_{l}=(-1)^{l-1}\sigma_{0}$, with $l$ labeling the layer index. Then the low-energy effective model for BTG can be derived as:
$H=\begin{pmatrix}h_{eff}(\lambda_{1})+h_{\Delta}(s_{1,1},t_{1,1})&h_{c}(s_{1,3},t_{1,3})\\ h_{c}(s_{1,3},t_{1,3})&h_{eff}(\lambda_{3})+h_{\Delta}(s_{3,3},t_{3,3})\end{pmatrix},$ (8)
@@
$h_{eff}(\lambda)=\begin{pmatrix}0&\lambda v_{3}\xi_{\mathbf{k}}-\frac{1}{\lambda\gamma_{1}}v^{2}_{0}(\xi_{\mathbf{k}}^{\dagger})^{2}\\ \lambda v_{3}\xi_{\mathbf{k}}^{\dagger}-\frac{1}{\lambda\gamma_{1}}v^{2}_{0}(\xi_{\mathbf{k}})^{2}&0\end{pmatrix},$ (9)
@@
$h_{\Delta}(s,t)=\begin{pmatrix}s\Delta+m&0\\ 0&t\Delta-m\end{pmatrix}$, and $h_{c}(s,t)=\begin{pmatrix}s\Delta&0\\ 0&t\Delta\end{pmatrix}$. The parameters in Eq. (8) can be found in Supplemental Material. The band structure near $K$ (or $K^{\prime}$) for the low-energy effective model of BTG with trigonal warping is shown in Fig. 3 (b). Here, the displacement field and staggered layer order are set to zero: $\Delta=0$ and $m=0$.
@@
Topological phase transitions in Bernal tetralayer graphene—The Hamiltonian in Eq. (8) can be interpreted as two coupled bilayer graphenes with different hopping parameters. In the following discussion, we refer to these two bilayer graphene systems as the first ( $H_{eff}^{1}=h_{eff}(\lambda_{1})+h_{\Delta}(s_{1,1},t_{1,1})$ ) and the second ( $H_{eff}^{2}=h_{eff}(\lambda_{3})+h_{\Delta}(s_{3,3},t_{3,3})$ ) bilayer graphene, respectively. Due to trigonal warping, the energy bands of both first and second bilayer graphene split into one central Dirac cone and three “leg” Dirac cones at the $K$ and $K^{\prime}$ valleys. In both bilayer graphenes, the central Dirac cone and the “leg” Dirac cone possess Berry phases of $-\zeta\pi$ and $\zeta\pi$, respectively, where $\zeta=\pm$ corresponds to the $K$ and $K^{\prime}$ valleys. Furthermore, although the central Dirac cones of these two bilayer graphenes overlap at the $K$ (or $K^{\prime}$) point, the momentum amplitudes $k$ of the three “leg” Dirac cones differ between these two systems, as shown in Fig. 3 (b). Specifically, since $\lambda_{3}>\lambda_{1}$, the momentum amplitude of the three “leg” Dirac cones in the second bilayer graphene ($k_{2}$) is larger than that in the first bilayer graphene ($k_{1}$). To highlight the effect of trigonal warping more clearly, the band structures along the $k_{y}=0$ line for BTG with $\Delta=0$ and $m=0$ are shown in Fig. 3 (c) and (d). Additionally, as shown in the low-energy effective model (Eq. (8)), the coupling strength between two bilayer graphenes is determined by the displacement field. For a fixed staggered layer order, the evolution of the band structure for the low-energy effective model of BTG with the increasing displacement field is illustrated in Fig. 4 (a).
The coupling between these two bilayer graphenes opens the gap at band-touching points and lowers the energy of the band near the central point of the $K$ (or $K^{\prime}$) valley, leading to a topological phase transition through a mass-inversion process. If mass inversion occurs in the bands belonging to the first bilayer graphene, the small momentum amplitude of the three “leg” Dirac cones causes the effective masses of both the central Dirac cone and the three “leg” Dirac cones to change sign simultaneously. In contrast, if mass inversion occurs in the bands belonging to the second bilayer graphene, only the effective mass of the central Dirac cone will change sign, as the three “leg” Dirac cones are located far from the central point of the $K$ (or $K^{\prime}$) valley.
@@
On the other hand, based on the effective masses of the first and second bilayer graphenes, we conclude that for $m>0$, when the interlayer potential satisfies $-3m<\Delta<0$, the amplitude of the effective mass for the first bilayer graphene, $M_{1}=|s_{1,1}\Delta+m|$, is smaller than that for the second bilayer graphene, $M_{2}=|s_{3,3}\Delta+m|$, i.e., $M_{1}<M_{2}$. As a result, the energy gap between the valence and conduction bands in the first bilayer graphene is smaller than that in the second bilayer graphene. Conversely, for $\Delta>0$, since $M_{1}>M_{2}$, the energy gap in the second bilayer graphene becomes smaller than that in the first bilayer graphene. The band structures corresponding to these two bilayer graphenes are shown in Fig. 4 (b). Therefore, for $m>0$, when $-3m<\Delta<0$, the topological phase transition results from mass inversion in the first bilayer graphene, which changes the sign of the effective mass at both the central Dirac cone and the three “leg” Dirac cones, causing the Chern number of the low-energy band to change by ±$\pm 2$. When $\Delta>0$, the phase transition arises from mass inversion in the second bilayer graphene, which only changes the sign of the effective mass at the central Dirac cone, leading to a change in the Chern number by ±$\pm 1$. For the case $m<0$, a similar result can be obtained: when $\Delta<0$ or $\Delta>-3m$, $M_{1}>M_{2}$; while for $0<\Delta<-3m$, $M_{1}<M_{2}$.
@@
Quantum anomalous Hall states in BTG—In analogy to the case in RPG, to investigate the QAH states in BTG, we assume that the staggered orders for the four flavors $K\uparrow$, $K\downarrow$, $K^{\prime}\uparrow$, and $K^{\prime}\downarrow$ are given by $H^{B}_{m}$, $H^{B}_{m}$, $H^{B}_{-m}$, and $H^{B}_{-m}$, respectively, with $m>0$. When $\Delta>3m$, the total Chern number in the $K$ valley for each spin is 3, while in the $K^{\prime}$ valley for each spin it is 0, resulting in a QAH state in BTG with a total Chern number of $C=6$. It is worth noting that for BTG, the topological phase transition behavior remains the same whether we consider the staggered layer or sublattice order. This is because the low-energy bands for BTG near the $K$ and $K^{\prime}$ points are primarily contributed by the sublattices $(A_{1},B_{2},A_{3},B_{4})$. When projecting the staggered layer and sublattice orders onto the basis formed by these sublattices, they yield the same expressions, leading to identical phase diagrams as the displacement field increases. Additionally, by using the full Hamiltonian of BTG [65, 72], we investigate the topological phase transitions. While the band structure differs from that of the low-energy effective model, the QAH state with a total Chern number $C=6$ persists (see Supplemental Material for details). Therefore, the low-energy effective model (Eq. (8)) provides a good description for the topological properties of the BTG.
@@
Conclusion and Discussion— In summary, we propose that for rhombohedral $N$-layer graphene, the emergence of QAH states with a Chern number $C\neq\pm N,\pm 2N$ can be attributed to an asynchronous mass inversion mechanism, which arises from the interplay between trigonal warping and staggered layer (or sublattice) order under a displacement field. The trigonal warping effect splits the low-energy bands of this multilayer graphene into the central touching point and “leg” Dirac cones in the $K$ (or $K^{\prime}$) valley. By incorporating the presence of staggered layer (or sublattice) order and increasing displacement fields, mass inversion can occur asynchronously for the central touching point and the “leg” Dirac cones, leading to a change in the Chern number that is not necessarily ±$\pm N$. Within this mechanism, we explain the appearance of the $C=-3$ QAH state in RPG which is not clear so far [21]. In addition, the topological phase transition observed in RPG with the increasing displacement field can be explained by assuming spin-polarized staggered sublattice and layer orders in the $K$ and $K^{\prime}$ valleys, respectively. Furthermore, under this assumption, as the displacement field increases from zero to sufficiently large values, the system transitions from a state with both spin and valley Chern numbers equal to zero to a quantum valley Hall state with a valley Chern number of 5, which is experimentally detectable. Moreover, it is worth noting that in experiments [21], QAH states are observed only under the application of a small, finite magnetic field. This observation may be attributed to the stabilization of valley polarization states, which arise from the coupling between orbital magnetization [73, 74, 75, 76, 77, 78, 79, 80] and the applied magnetic field. Interestingly, the asynchronous mass inversion mechanism can also be applied to BMG, revealing the emergence of the $C=6$ QAH state in BTG, which can be observed in experiments.
@@
Acknowledgements—K. T. L. acknowledges the support of the Ministry of Science and Technology, China, and Hong Kong Research Grant Council through Grants No. 2020YFA0309600, No. RFS2021-6S03, No. C6025-19G, No. AoE/P-701/20, No. 16310520, No. 16307622, and No. 16309223.
SUPPLEMENTAL MATERIAL NOTE 1: RHOMBOHEDRAL MULTILAYER GRAPHENE
.1 Low-energy effective model for rhombohedral multilayer graphene
@@
where $N$ denotes the number of layers. Therefore, the matrix form of the Hamiltonian for RMG is written as:
@@
where $U_{l}=(-1)^{l-1}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$, $O=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$, $l$ is the layer index, and $m$ is the amplitude of this order. $Diag(...)$ denotes a diagonal matrix whose diagonal elements are listed in the brackets. In addition, the Hamiltonian for interlayer potential induced by displacement field can be expressed as:
@@
In this effective mass (Eq. (S6)), $m$ is the amplitude of the staggered layer order, $\Delta$ is the amplitude of the interlayer potential induced by the displacement field, $a_{0}$ is the lattice constant of the graphene, and $k^{2}=\xi_{\mathbf{k}}\xi^{\dagger}_{\mathbf{k}}=|k_{x}+ik_{y}|^{2}$. As discussed in the main text, the combination of trigonal warping and the momentum dependence of the effective mass is responsible for the appearance of additional quantum anomalous Hall (QAH) states with Chern number $C\neq\pm N,\pm 2N$ in rhombohedral $N$-layer graphene. The effective hopping term $X_{N}(\mathbf{k})$ for two-band model is written as [50]:
@@
In rhombohedral pentalayer graphene, the band structure of the low-energy effective model under different displacement fields and a fixed staggered layer order is shown in Fig.2 of the main text. As the displacement field increases, the effective mass term for the three “leg” Dirac cones first changes sign at $\Delta=-m_{I}<-m$ (with $m<0$), resulting in a topological phase transition with a change in the Chern number of ±$\pm 3$. As the displacement field continues to increase, the effective mass for the central band-touching point (quadratic touching) also changes sign at $\Delta=-m$, leading to a second topological phase transition with a change in the total Chern number of ±$\pm 2$.
Figure S2: (a). The evolution of the band structure for a fixed staggered layer order (SLO) $m=-10meV$ with increasing $\Delta$. (b). The evolution of the band structure for a fixed staggered sublattice order (SSO) $m=-15meV$ with increasing $\Delta$. Other parameters in (a) and (b) are: $v_{0}=\frac{\sqrt{3}}{2}\times 3160~{}meV$, $\gamma_{1}=390~{}meV$, $\gamma_{2}=-20~{}meV$, $v_{3}=\frac{\sqrt{3}}{2}\times 315~{}meV$, and $v_{4}=\frac{\sqrt{3}}{2}\times 40~{}meV$.
.2 The band structure and topological phase transition for the full Hamiltonian of RPG
In this section, to make the analysis more practical, we calculate the band structure of RPG using the full Hamiltonian shown in Eq. (S2) instead of the low-energy effective model, accounting for the staggered layer or sublattice order (SLO or SSO) and displacement field. The evolution of the bands with increasing displacement field for a fixed SLO and SSO, are shown in Fig. S2 (a) and (b), respectively. Although the band structure differs from the low-energy two-band model, the two-step topological phase transition behavior remains unchanged. Therefore, we can conclude that the low-energy two-band model can capture the key physical mechanism underlying the QAH states in RPG.
.3 The effect of trigonal warping term $\gamma_{3}$ in rhombohedral pentalayer graphene
In this section, we discuss the effect of the trigonal warping term $\gamma_{3}$ on the low-energy effective model for rhombohedral pentalayer graphene (RPG). By using the perturbation theory mentioned in Sec. .1, the two-band effective Hamiltonian for RPG can be written as:

### v2

@@ Abstract
Figure 3: (a). Side view of the Bernal tetralayer graphene (BTG), the blue and red atoms represent the $A$ and $B$ sublattices, respectively. The hopping parameters are indicated by different colors ($v_{i}=\frac{\sqrt{3}}{2}\gamma_{i}$): $\gamma_{0}=3000~{}meV$, $\gamma_{1}=400~{}meV$, $\gamma_{2}=-20~{}meV$, $\gamma_{3}=300~{}meV$, and $\gamma_{4}=\gamma_{5}=40~{}meV$ Koshino and McCann (2011). Here, the displacement field and staggered layer orders are set to zero: $\Delta=0$, and $m=0$. (b). The band structure near $K$ (or $K^{\prime}$) valley in the presence of the trigonal warping. The bands in the green circle show the enlarged view near $K$ (or $K^{\prime}$) point. (c). The band structure for the low-energy effective model along $k_{y}=0$ line with and without $\gamma_{3}$, which are shown by red solid and blue dashed lines, respectively. Here, we set $\Delta=0$ and $m=0$. (d). Enlarged view of the band structure for the region marked by the green circle in (c).
Furthermore, to make the analysis more practical, we investigate the band structure and topological phase transition using the full Hamiltonian of RPG (Eq. 1), taking into account the staggered layer order and displacement field. Although the additional trigonal warping term, $\gamma_{3}$, causes a slight further splitting at the central band touching point in the $K$ (or $K^{\prime}$) valley, and the interlayer hopping term, $\gamma_{4}$, introduces particle-hole asymmetry, the asynchronous mass inversion mechanism and the phase diagram for RPG remain unaffected. The results of the full model calculations are shown in Supplemental Material sup .
Generalizing to other rhombohedral multilayer graphene systems—To verify the universality of the asynchronous mass inversion mechanism in multilayer graphene systems and pave the way for experimental verification, we derive the low-energy effective model for rhombohedral $N$-layer graphene ($N\geq 3$) in the presence of staggered layer order and displacement field (see Supplemental Material sup ). By applying this model to rhombohedral tetralayer graphene (RTG), we can get the effective mass for RTG ($m<0$):
$M_{4,\text{eff}}(k)=(m+\Delta)-\frac{(3m-\Delta)v^{2}a_{0}^{2}k^{2}}{3\gamma_{1}^{2}}.$ (7)
Thus, the mass-inversion conditions for the central Dirac cone and the satellite Dirac cones are $\Delta=\Delta_{1}=-m$ and $\Delta=\Delta_{2}=\frac{3\epsilon^{\prime}-1}{1+\epsilon^{\prime}}m$, respectively. Under the condition $0<\epsilon^{\prime}=\frac{v_{0}^{2}a_{0}^{2}k^{2}}{3\gamma_{1}^{2}}<\frac{1}{3}$, we derive that $\Delta_{2}<\Delta_{1}$, which predicts the emergence of a $C=3$ QAH state prior to the experimentally observed $C=4$ QAH state as the displacement field increases in RTG.
The effective model for Bernal tetralayer graphene—Next, we apply the asynchronous mass inversion mechanism to explain the recently observed $C=6$ QAH state in BTG Cho et al. (2025). The side view and hopping parameters for BTG are shown in Fig. 3 (a), and the staggered layer order is written as:
@@
where $U_{l}=(-1)^{l-1}\sigma_{0}$, with $l$ labeling the layer index. Then the low-energy effective model for BTG can be derived as Koshino and McCann (2011):
$H=\begin{pmatrix}h_{\text{eff}}(\lambda_{1})+h_{\Delta}(s_{1,1},t_{1,1})&h_{c}(s_{1,3},t_{1,3})\\ h_{c}(s_{1,3},t_{1,3})&h_{\text{eff}}(\lambda_{3})+h_{\Delta}(s_{3,3},t_{3,3})\end{pmatrix},$ (9)
@@
$h_{eff}(\lambda)=\begin{pmatrix}0&\lambda v_{3}\xi_{\mathbf{k}}-\frac{1}{\lambda\gamma_{1}}v^{2}_{0}(\xi_{\mathbf{k}}^{\dagger})^{2}\\ \lambda v_{3}\xi_{\mathbf{k}}^{\dagger}-\frac{1}{\lambda\gamma_{1}}v^{2}_{0}(\xi_{\mathbf{k}})^{2}&0\end{pmatrix},$
@@
$h_{\Delta}(s,t)=\begin{pmatrix}s\Delta+m&0\\ 0&t\Delta-m\end{pmatrix}$, and $h_{c}(s,t)=\begin{pmatrix}s\Delta&0\\ 0&t\Delta\end{pmatrix}$. The parameters in Eq. (9) are $\lambda_{1}=(\sqrt{5}-1)/2$, $\lambda_{3}=(\sqrt{5}+1)/2$, $s_{1,1}=-t_{1,1}=\frac{1}{15}(5+2\sqrt{5})$, $s_{3,3}=-t_{3,3}=\frac{1}{15}(5-2\sqrt{5})$, and $s_{1,3}=t_{1,3}=\frac{4}{3\sqrt{5}}$. The band structure near $K$ (or $K^{\prime}$) point for the low-energy effective model of BTG with trigonal warping is shown in Fig. 3 (b). In this figure, the displacement field and staggered layer order are set to zero: $\Delta=0$ and $m=0$.
@@
Topological phase transitions in Bernal tetralayer graphene—The Hamiltonian in Eq. (9) can be interpreted as two coupled bilayer graphenes with different hopping parameters. In the following discussion, we refer to these two bilayer graphene systems as the first ( $H_{\text{eff}}^{1}=h_{\text{eff}}(\lambda_{1})+h_{\Delta}(s_{1,1},t_{1,1})$ ) and the second ( $H_{\text{eff}}^{2}=h_{\text{eff}}(\lambda_{3})+h_{\Delta}(s_{3,3},t_{3,3})$ ) bilayer graphene, respectively. Due to the trigonal warping, the energy bands of both the first and second bilayer graphene split into one central Dirac cone and three satellite Dirac cones at the $K$ and $K^{\prime}$ valleys. The central Dirac cone and the satellite Dirac cone possess Berry phases of $-\zeta\pi$ and $\zeta\pi$, respectively, where $\zeta=\pm$ corresponds to the $K$ and $K^{\prime}$ valleys. Moreover, since $\lambda_{3}>\lambda_{1}$, the momentum amplitude of the three satellite Dirac cones in the second bilayer graphene ($k_{2}$) is larger than that in the first bilayer graphene ($k_{1}$), as shown in Fig. 3 (b). Fig. 3 (c) and (d) show the band structures along the $k_{y}=0$ line for BTG with $\Delta=0$ and $m=0$, explicitly demonstrating the trigonal warping effect. Additionally, as shown in the low-energy effective model (Eq. (9)), the coupling strength between the two bilayer graphenes ($h_{c}$) is determined by the displacement field. For a fixed staggered layer order, the evolution of the band structure for BTG with the increasing displacement field is illustrated in Fig. 4 (a).
@@
We next analyze the mass inversion process as the displacement field increases in BTG. Based on the effective masses ($h_{\Delta}$) of the first and second bilayer graphenes (in Eq. 9), we conclude that for $m>0$, when the interlayer potential satisfies $-3m<\Delta<0$, the amplitude of the effective mass for the first bilayer graphene, $M_{1}=|s_{1,1}\Delta+m|$, is smaller than that for the second bilayer graphene, $M_{2}=|s_{3,3}\Delta+m|$, i.e., $M_{1}<M_{2}$. As a result, the energy gap between the valence and conduction bands in the first bilayer graphene is smaller than that in the second bilayer graphene. Consequently, for $m>0$, when $-3m<\Delta<0$, the topological phase transition arises from mass inversion in the band whose component is from the first bilayer graphene. Due to the small momentum amplitude of the three satellite Dirac cones in the first bilayer graphene, the effective masses of the central Dirac cone and the satellite Dirac cones invert simultaneously, resulting in a change of ±$\pm 2$ in the Chern number of the low-energy valence band. Conversely, for $\Delta>0$, since $M_{1}>M_{2}$, the energy gap in the second bilayer graphene becomes smaller than that in the first bilayer graphene. Therefore, when $\Delta>0$, the topological phase transition arises from mass inversion in the band whose component is from the second bilayer graphene. Due to the large momentum amplitude of the satellite Dirac cones in the second bilayer graphene, the mass inversion only occurs at the central Dirac cone. This results in a change of ±$\pm 1$ in the Chern number. The band structures corresponding to two bilayer graphenes with $m>0$ are shown in Fig. 4 (b). On the other hand, for the case $m<0$, a similar result can be obtained: when $\Delta<0$ or $\Delta>-3m$, $M_{1}>M_{2}$, the Chern number changes by ±$\pm 1$; while for $0<\Delta<-3m$, $M_{1}<M_{2}$, the Chern number changes by ±$\pm 2$.
@@
Quantum anomalous Hall states in BTG—To investigate the QAH states in BTG, we assume that the staggered orders for the four flavors $K\uparrow$, $K\downarrow$, $K^{\prime}\uparrow$, and $K^{\prime}\downarrow$ are given by $H^{B}_{m}$, $H^{B}_{m}$, $H^{B}_{-m}$, and $H^{B}_{-m}$, respectively, with $m>0$. When $\Delta>3m$, the total Chern number of valence bands in the $K$ valley for each spin is 3, while in the $K^{\prime}$ valley for each spin, it is 0. This results in a QAH state in BTG with a total Chern number of $C=6$. Additionally, by using the full Hamiltonian of BTG Ghazaryan et al. (2023); Chen et al. (2023), we investigate its band structure and topological phase transitions. While the band structure differs from that of the low-energy effective model and the topological phase transition varies with different amplitudes of the staggered layer order $m$, the QAH state with a total Chern number of $C=6$ persists (see Supplemental Material for details sup ).
@@
Conclusion and Discussion— In summary, we propose that for rhombohedral $N$-layer graphene, the emergence of QAH states with a Chern number $C\neq\pm N,\pm 2N$ can be attributed to an asynchronous mass inversion mechanism, which arises from the interplay between trigonal warping, staggered layer order, and displacement field. The trigonal warping splits the low-energy bands of this multilayer graphene into the central touching point and satellite Dirac cones in the $K$ (or $K^{\prime}$) valley. With staggered layer order and increasing displacement fields, mass inversion can occur asynchronously at the central touching point and the satellite Dirac cones, leading to Chern number changes that are not necessarily equal to ±$\pm N$. Within this mechanism, we explain the origin of the $C=-3$ QAH state in RPG Han et al. (2024a). In addition, we predict the appearance of the QAH state with $C=3$ in RTG, which is experimentally detectable. Interestingly, the asynchronous mass inversion mechanism can also be applied to BTG, revealing the emergence of the $C=6$ QAH state, which has been observed experimentally Cho et al. (2025). Moreover, it is worth noting that in experiments Han et al. (2024a); Cho et al. (2025), QAH states are observed only under the application of a small, finite magnetic field. This observation may be attributed to the stabilization of valley-polarized orders, resulting from the coupling between orbital magnetization Thonhauser et al. (2005); Xiao et al. (2005); Ceresoli et al. (2006); Shi et al. (2007); Xiao et al. (2010); Bianco and Resta (2013, 2016); Zhu et al. (2020) and the applied magnetic field. This coupling remains to be explored in future studies.
@@
Acknowledgements—The authors thank Long Ju and Fan Zhang for inspiring discussions. K. T. L. acknowledges the support of the Ministry of Science and Technology, China, and Hong Kong Research Grant Council through Grants No. 2020YFA0309600, No. RFS2021-6S03, No. C6025-19G, No. AoE/P-701/20, No. 16310520, No. 16307622, and No. 16309223.
SUPPLEMENTAL MATERIAL NOTE I: RHOMBOHEDRAL MULTILAYER GRAPHENE
.1 I-A. Low-energy effective model for rhombohedral multilayer graphene ($N\geq 3$)
@@
where $A_{i}$ and $B_{i}$ represent the $A$ and $B$ sublattice in the $i$ layer, and $N$ denotes the number of layers (with the assumption that $N\geq 3$). Therefore, the matrix form of the Hamiltonian for RMG is written as:
@@
where $U_{l}=(-1)^{l-1}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$, $O=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}$, $l$ is the layer index, and $m$ is the amplitude of this order. $Diag(...)$ denotes a diagonal matrix whose diagonal elements are listed in the brackets. In addition, the Hamiltonian for interlayer potential induced by the displacement field can be expressed as:
@@
In this effective mass (Eq. (S6)), $m$ is the amplitude of the staggered layer order, $\Delta$ is the amplitude of the interlayer potential induced by the displacement field, $a_{0}$ is the lattice constant of the graphene, and $k^{2}=\xi_{\mathbf{k}}\xi^{\dagger}_{\mathbf{k}}=|k_{x}+ik_{y}|^{2}$. As discussed in the main text, the combination of trigonal warping and the momentum dependence of the effective mass is responsible for the appearance of additional quantum anomalous Hall (QAH) states with Chern number $C\neq\pm N,\pm 2N$ in rhombohedral $N$-layer graphene. The effective hopping term $X_{N}(\mathbf{k})$ for the two-band model is written as Koshino and McCann (2009):
@@
In rhombohedral pentalayer graphene (RPG), the band structure of the low-energy effective model under different displacement fields and a fixed staggered layer order is shown in Fig.2 of the main text. As the displacement field increases, the effective mass term for the three satellite Dirac cones first changes sign at $\Delta=-m_{I}<-m$ (with $m<0$), resulting in a topological phase transition with a change in the Chern number of ±$\pm 3$. As the displacement field continues to increase, the effective mass for the central band-touching point (quadratic touching) also changes sign at $\Delta=-m$, leading to a second topological phase transition with a change in the total Chern number of ±$\pm 2$.
.2 I-B. The band structure and topological phase transition for the full Hamiltonian of RPG
In this section, to make the analysis more practical, we calculate the band structure of RPG using the full Hamiltonian shown in Eq. (S2) instead of the low-energy effective model, accounting for the staggered layer order and displacement field. The evolution of the bands with increasing displacement field for a fixed staggered layer order is shown in Fig. S2. Although the band structure differs from the low-energy two-band model, the asynchronous mass inversion behavior remains unchanged. Therefore, the low-energy two-band model can capture the key physical mechanism underlying the formation of the QAH states in RPG.
Figure S2: The evolution of the band structure for a fixed staggered layer order $m=-10meV$ with increasing $\Delta$. Other parameters in are: $v_{0}=\frac{\sqrt{3}}{2}\times 3160~{}meV$, $\gamma_{1}=390~{}meV$, $\gamma_{2}=-20~{}meV$, $v_{3}=\frac{\sqrt{3}}{2}\times 315~{}meV$, and $v_{4}=\frac{\sqrt{3}}{2}\times 40~{}meV$.
.3 I-C. The effect of trigonal warping term $\gamma_{3}$ in rhombohedral pentalayer graphene
In this section, we examine the effect of the trigonal warping term $\gamma_{3}$ in rhombohedral pentalayer graphene (RPG). By using the perturbation theory mentioned in Sec.I-A, the two-band effective Hamiltonian for RPG can be written as: